# Metastable Strings and Gravitational Waves in One-Scale Models

### v1

@@ Abstract
Metastable cosmic strings provide a minimal and predictive origin for the stochastic gravitational-wave background reported by Pulsar Timing Array experiments. We analyse this possibility in electroweak-like dark sectors with a single-stage breaking $SU(2)\times U(1)\!\to\!U(1)$ driven by one Higgs field. In the regime with dark sector Higgs mass below the $Z^{\prime}$ mass, and for sufficiently small $W^{\prime}$ mass, the resulting $Z$-string is classically stable but undergoes quantum decay via nucleation of monopole–antimonopole pairs along the string. We compute the corresponding semiclassical bounce action in a thin-defect approximation, treating both the string core and the monopole endpoints as localised defects whose sizes are small relative to their separation in the tunnelling configuration. This yields a decay rate per unit length that depends on the gauge couplings and the mass hierarchy. We delineate the parameter space in which single-scale dark-sector models reproduce the PTA signal, and verify that the thin-defect approximation remains valid throughout the phenomenologically favoured region, without invoking extended Higgs sectors or multi-stage symmetry breaking.
@@ 1 Introduction
Gravitational wave (GW) astronomy has entered a new era with the recent discoveries by Pulsar Timing Array (PTA) collaborations. In 2023, NANOGrav [NANOGrav:2023gor, NANOGrav:2023icp, NANOGrav:2023hfp, NANOGrav:2023ctt, NANOGrav:2023hvm], EPTA [EPTA:2023fyk, EPTA:2023sfo, EPTA:2023akd, EPTA:2023gyr, EPTA:2023xxk, EuropeanPulsarTimingArray:2023egv], PPTA [Zic:2023gta, Reardon:2023zen, Reardon:2023gzh], and CPTA [Xu:2023wog] all reported compelling evidence for a stochastic GW background in the nanohertz frequency range. The detection of quadrupolar spatial correlations consistent with the Hellings–Downs curve [Hellings:1983fr] strongly indicates that the signal is of gravitational origin. While the most anticipated source is a cosmic population of inspiralling supermassive black hole binaries [Haehnelt:1994wt, Rajagopal:1994zj, Jaffe:2002rt, Wyithe:2002ep, Sesana:2004sp, Sesana:2008mz, Burke-Spolaor:2018bvk, Middleton:2020asl], the origin is not yet confirmed, and several cosmological explanations remain plausible. These include first-order phase transitions [NANOGrav:2021flc, Kamionkowski_1994, Grojean:2006bp, Huber:2008hg, Jaeckel_2016, Ellis:2018mja, Nakai:2020oit, Ratzinger:2020koh, Ashoorioon:2022raz, Freese:2022qrl, Morgante:2022zvc, Bringmann:2023opz, Winkler:2024olr], annihilating domain walls [Ferreira:2022zzo], cosmic strings [Blasi:2020mfx, Ellis:2020ena, Buchmuller:2020lbh, Samanta:2020cdk, King:2020hyd, Lazarides:2021uxv, Buchmuller:2023aus, Chitose:2023dam], axion dynamics [Ratzinger:2020koh, Wang:2022rjz], and GWs induced by large scalar fluctuations [Vaskonen:2020lbd, Kohri:2020qqd, DeLuca:2020agl, Zhao:2022kvz, Dandoy:2023jot]. The PTA era therefore offers a unique opportunity to probe physics far beyond the reach of terrestrial experiments.
@@
In this work we explore the possibility that the PTA signal originates from metastable cosmic strings. Unlike their stable counterparts, metastable strings decay over time, leading to a distinct, time-evolving GW spectrum. Their lifetime is determined by a quantum tunnelling process in which a segment of string nucleates a monopole–antimonopole pair that severs it [Preskill:1992ck]. The decay rate is exponentially suppressed by a semiclassical bounce action, conventionally parametrised by a dimensionless quantity $\kappa$, whose magnitude controls both the amplitude and spectral slope of the resulting GW background. Current PTA data already prefer $\kappa$ values within a relatively narrow range [NANOGrav:2023hvm, Buchmuller:2023aus, Antusch:2023zjk, Madge:2023dxc], motivating microscopic models that can realise this parameter naturally. We focus on a simple and predictive setting capable of generating such metastable strings: an electroweak-like dark-sector gauge theory with a single-stage symmetry breaking, $SU(2)\times U(1)\ \to\ U(1),$ which can be viewed as a minimal analogue of the Standard Model’s electroweak sector. In this theory, the resulting $Z$-strings are not topologically protected but can be dynamically long-lived in the near-semi-local regime, where the $SU(2)$ coupling is small compared to the $U(1)$ one. Their decay proceeds through monopole–antimonopole nucleation, in contrast to multi-stage models ($SU(2)\!\to\!U(1)\!\to\!\emptyset$) where monopoles are topologically embedded from the outset [Buchmuller:2019gfy, Buchmuller:2023aus, Shifman:2002yi, Chitose:2023dam, Chitose:2025cmt].
@@ 2.1 Lagrangian, symmetry breaking and mass eigenstates
A key feature of our setup is that it involves only a single stage of symmetry breaking, governed by one vev $\eta$, as shown in Eq. (2.1) [^1] This is different from multi-scale models of metastable strings, where the symmetry breaking pattern is governed by different vevs, that are often considered in the literature, see e.g. [Buchmuller:2019gfy, Buchmuller:2023aus, Shifman:2002yi, Chitose:2023dam, Chitose:2025cmt].. The Lagrangian of this theory is
@@ 2.2 String ansatz and boundary conditions
The model described above admits classical solutions that are static and uniform in one spatial direction [Vachaspati:1992fi]. They are characterised by a pair of field configurations, $Z^{\,\rm str}_{\mu}(x_{1},x_{2})$ and $\Phi^{\rm str}(x_{1},x_{2})$, that have no time ($t$) and $z$ ($x_{3}$) dependence and correspond to a string sourced by the $Z$ and Higgs fields along the $x_{3}$-direction:
@@
The field configurations of Eqs. (2.10) and (2.11) are known as the electroweak strings, or $Z$-strings, and amount to a non-Abelian embedding of the Abrikosov-Nielsen-Olesen (ANO) classical string of the Abelian Higgs model [Abrikosov:1956sx, Nielsen:1973cs]. The index $n\in\mathbb{Z}$ is the magnetic flux of the $Z$-string, and in what follows, we will restrict to an elementary string with $n=1$. Generally, the field configurations described by the ansatz of Eqs. (2.10) and (2.11) are not guaranteed or expected to be stable throughout the model’s parameter space.
@@ 2.3 Topology and the semi-local limit
Although not protected by topology, Refs. [Hindmarsh:1991jq, James:1992wb] showed that such electroweak-like string configurations can be classically stable local minima [^2] A note on terminology: A field configuration is called classically stable if it is a local minimum of the energy functional $E$. In particular, there should be no negative or zero modes when expanding $E$ around a classically stable field configuration. There can, and in general will exist other local minima with lower values of $E$ separated by finite or infinite energy barriers from the classical solution. When the barrier is finite, the classical string can decay via quantum tunnelling into the true minimum. In this case the classical string is metastable and, in favourable conditions, its quantum decay rate can be computed semiclassically [Preskill:1992ck] using a version of Coleman’s false vacuum decay approach [Coleman:1977py]. of the energy per unit length in a restricted region of parameter space when the model becomes semi-local:
@@
We demonstrate this stability in the semi-local limit, where the $SU(2)$ sector of the gauge group $G$ becomes a global symmetry, while the $U(1)$ factor remains gauged. The resulting semi-local model then possesses a topologically conserved, integer-valued magnetic flux [Vachaspati:1991dz, Preskill:1992bf] (see also sec. IID.2 in Ref. [Preskill:1992ck] for a concise summary of the topological argument), which corresponds to the quantum number $n\in\mathbb{Z}$ appearing in the string field configurations Eqs. (2.10) and (2.11). As mentioned, in what follows, we will restrict to the elementary flux sector with $n=1$. Stability is determined by an energy comparison between two configurations: the localised $Z$-string, where energy is confined in a flux tube of tension $\mu_{\rm str}$, and a delocalised texture, where the same flux spreads across a broad scalar field configuration of tension $\mu_{\infty}$:
@@ 3 Tension Parameters and Classical Stability
This section computes the two dimensionless tension coefficients that control the $Z$-string’s stability and decay rate. In Section 3.1 we derive $\alpha_{\infty}$ (delocalised flux, sigma-model limit); in Section 3.2 we compute the string tension $\alpha_{\rm str}(\beta)$ numerically as a function of the mass ratio $\beta$
@@ 3.1 Computing the $\alpha_{\infty}$ tension coefficient
To compute $\alpha_{\infty}$, we evaluate the energy of static two-dimensional configurations carrying one unit of magnetic flux, uniformly delocalised over a disc of radius $R$ and taking the $R\to\infty$ limit. Following Ref. [Preskill:1992ck] (sec. VIIIA), we can make two simplifying assumptions. For these infinitely-spread finite energy configurations, the Higgs field takes values in the vacuum manifold everywhere,
@@
where $g(\Phi)_{\alpha\beta}$ is the metric tensor of the $CP^{1}$ theory in affine coordinates, given by the Fubini-Study metric, see $e.g.$ [Eguchi:1980jx, Hindmarsh:1992yy],
@@
The real-valued fields $\varphi^{a}$ defined in Eq. (3.6) satisfy the relation $(\varphi^{1})^{2}+(\varphi^{1})^{2}+(\varphi^{1})^{2}=\eta^{2}/2$ and conveniently parameterise the vacuum manifold of our original Abelian Higgs model,
@@
Here $E$ is the transverse energy, $i.e.$ the 3+13{+}1D string tension (energy per unit length). The model of Eq. (3.9) has topological solitons [Polyakov:1975yp, Rubakov:2002fi] with topological charges $n\in\pi_{2}(S^{2})\sim\mathbb{Z}$. The single soliton solution is given by,
@@
More detail and a derivation of the soliton solution above can also be found in Sec. 7 of the Ref. [Rubakov:2002fi]. In the sigma model of Eq. (3.9), the soliton size parameter $r_{0}$ is a flat direction. But the sigma model soliton provides an accurate approximation to the large-size solution of our $SU(2)\times U(1)$ semi-local model only for large scale sizes. Hence, we can identify $r_{0}$ with $R$ when we take the latter to infinity. By equating the two-dimensional soliton mass in Eq. (3.11) with the string tension $\mu_{\infty}$ of the infinite-size magnetic flux configuration of Eq. (2.16), we conclude that
@@ 3.2 Computing the string tension coefficient
The dependence of $\alpha_{\rm str}$ on the mass ratio $\beta$ is displayed in the right panel of Figure 1. As expected from the behaviour of the profile functions, $\alpha_{\rm str}$ increases slowly as $\beta$ increases: a heavier Higgs field localises the scalar core and slightly raises the overall tension. The role of the $SU(2)$ gauge coupling $g$ is implicit in the choice of dimensionless variables and does not appear in the energy functional [Vachaspati:1992fi], or alter the shape of the curves shown in the right panel of Figure 1, although it re-enters when converting back to physical units.
Figure 1: Left: Profile functions P(r~)P($\tilde${r}) and V(r~)V($\tilde${r}) for the classical string configuration. The functions are plotted for $\beta=0.02,\,0.2$ and 1. These functions are independent of $g$. Right: The blue line shows the dependence of $\alpha_{\text{str}}$ on $\log\beta$, for the range $2<\log_{10}\beta<0$, where the string solution can be shown to be metastable. Note also that $\alpha_{\text{str}}$ depends on $\beta$ only, and has no further dependence on $g$. The grey dotted line shows the analytically approximation for $\alpha_{\rm str}$ valid in the small-$\beta$ limit and the red line $\alpha_{\rm str}=\alpha_{\infty}=\pi$.
@@
We observe that for $\beta=1$ the string coupling satisfies $\alpha_{\rm str}=\pi$, coinciding with the value for $\alpha_{\infty}$ and saturating the Bogomolnyi bound [Bogomolny:1975de] for the energy. In addition, in the $\beta\to 0$ limit the ANO string tension has been obtained analytically in Ref. [Shifman:2002yi],
@@ 4 Bounce Calculation
When we turn on $g\neq 0$ the $SU(2)$ group is gauged and the string magnetic flux is no longer protected by topology as it can ‘unwind’ by exciting non-Abelian gauge field degrees of freedom. The string is no longer absolutely stable but there is a region in the parameter space where it is still a local minimum now separated by a finite energy barrier from minima with lower energy. This classical stability region of the string was established in Ref. [James:1992wb], and is shown in Figure 3 as the area to the right of the black contour on the $(\beta,\sin^{2}\theta_{w})$ plane. The classical string decays by quantum tunnelling and in this section, we will derive the bounce action which determines the semiclassical exponent of the string decay rate following the thin-defect aka thin-wall approach of [Preskill:1992ck].
@@
can terminate, while the magnetic flux of the $B_{\mu}$ component is converted to the magnetic flux of the massless photon field $A_{\mu}$. The end points of the $Z$-string are electromagnetic monopoles of Nambu [Nambu:1977ag, Kibble:2015twa], [^3] These configurations exist only as string endpoints, i.e. they are always confined by the string and do not appear as free classical solutions. In practice they are ’t Hooft–Polyakov-type monopole configurations composed of the $W_{\mu}^{a}$ gauge field and the adjoint Higgs field $n^{a}=\Phi^{\dagger}\tau^{a}\Phi/(\Phi^{\dagger}\Phi)$. This composite scalar field configuration is not invertible for $\Phi(x)=0$, which is precisely the point where the $Z$-string is attached. they absorb the string’s $Z$-flux and continuously convert it into $A$-flux,
@@
and considering the hierarchy of scales shown in the left panel of Figure 2. Summing these three contributions gives the total energy contained within the monopole core [Preskill:1992ck]:
@@
where again we expanded in $g^{2}/g^{\prime 2}\ll 1$ [^4] We note that in [Preskill:1992ck], there is an error in the calculation of $R_{\rm core}^{2}$ by a factor of 2.. To determine the bounce action, we follow the Euclidean picture shown in the right panel of Figure 2, reproducing the monopole–antimonopole nucleation process on aă string as described Ref. [Preskill:1992ck]. A string sweeps out a two‑dimensional worldsheet with one timelike and one spacelike coordinate. Performing a Wick rotation converts the timelike direction into another spacelike one, converting the Lorentzian worldheet into a Euclidean surface. Quantum fluctuations may then punch a circular hole of radius $R$ in the worldsheet and the rim of this hole is the closed worldline of the nucleated monopole–antimonopole pair. In the thin‑wall limit, where the string thickness and the monopole core size are negligible compared with the radius of the critical bubble $R$, the Euclidean action simplifies to
@@
The result is valid as long as all scales remain much smaller than $R$. The monopole mass in this expression corresponds to the energy contained within the monopole core, obtained in Eq. (4.7), and combining these two equations we find, reproducing the result found in Ref. [Preskill:1992ck],

### v2

@@ Abstract
Metastable cosmic strings provide a minimal and predictive origin for the stochastic gravitational-wave background reported by Pulsar Timing Array experiments. We analyse this possibility in electroweak-like dark sectors with a single-stage breaking $SU(2)\times U(1)\!\to\!U(1)$ driven by one Higgs field. In the regime with dark sector Higgs mass below the $Z^{\prime}$ mass, and for sufficiently small $W^{\prime}$ mass, the resulting $Z$-string is classically stable but undergoes quantum decay via nucleation of monopole–antimonopole pairs along the string. We compute the corresponding semiclassical bounce action in a thin-defect approximation, treating both the string core and the monopole endpoints as localised defects whose sizes are small relative to their separation in the tunnelling configuration. This yields a decay rate per unit length that depends on the gauge couplings and the mass hierarchy. We delineate the parameter space in which single-scale dark-sector models reproduce the PTA signal, demonstrating the applicability of the thin-defect approximation throughout the phenomenologically favoured region and without invoking extended Higgs sectors or multi-stage symmetry breaking.
@@ 1 Introduction
Gravitational wave (GW) astronomy has entered a new era with the recent discoveries by Pulsar Timing Array (PTA) collaborations. In 2023, NANOGrav [7, 47, 5, 6, 4], EPTA [10, 8, 9, 12, 11, 76], PPTA [87, 69, 68], and CPTA [84] all reported compelling evidence for a stochastic GW background in the nanohertz frequency range. The detection of quadrupolar spatial correlations consistent with the Hellings–Downs curve [40] strongly indicates that the signal is of gravitational origin. While the most anticipated source is a cosmic population of inspiralling supermassive black hole binaries [39, 66, 45, 83, 73, 74, 23, 56], the origin is not yet confirmed, and several cosmological explanations remain plausible. These include first-order phase transitions [14, 48, 38, 43, 44, 33, 60, 67, 15, 36, 59, 18, 82], annihilating domain walls [35], cosmic strings [16, 34, 21, 72, 50, 52, 22, 25], axion dynamics [67, 81], and GWs induced by large scalar fluctuations [80, 51, 29, 86, 28]. The PTA era therefore offers a unique opportunity to probe physics far beyond the reach of terrestrial experiments.
@@
In this work we explore the possibility that the PTA signal originates from metastable cosmic strings. Unlike their stable counterparts, metastable strings decay over time, leading to a distinct, time-evolving GW spectrum. Their lifetime is determined by a quantum tunnelling process in which a segment of string nucleates a monopole–antimonopole pair that severs it [64]. The decay rate is exponentially suppressed by a semiclassical bounce action, conventionally parametrised by a dimensionless quantity $\kappa$, whose magnitude controls both the amplitude and spectral slope of the resulting GW background. Current PTA data already prefer $\kappa$ values within a relatively narrow range [4, 22, 13, 54], motivating microscopic models that can realise this parameter naturally. We focus on a simple and predictive setting capable of generating such metastable strings: an electroweak-like dark-sector gauge theory with a single-stage symmetry breaking, $SU(2)\times U(1)\ \to\ U(1),$ which can be viewed as a minimal analogue of the Standard Model’s electroweak sector. In this theory, the resulting $Z$-strings are not topologically protected but can be dynamically long-lived in the near-semi-local regime, where the $SU(2)$ coupling is small compared to the $U(1)$ one. Their decay proceeds through monopole–antimonopole nucleation, in contrast to multi-stage models ($SU(2)\!\to\!U(1)\!\to\!\emptyset$) where monopoles are topologically embedded from the outset [20, 22, 75, 25, 26].
@@ 2.1 Lagrangian, symmetry breaking and mass eigenstates
A key feature of our setup is that it involves only a single stage of symmetry breaking, governed by one vev $\eta$, as shown in Eq. (2.1) [^1] This is different from multi-scale models of metastable strings, where the symmetry breaking pattern is governed by different vevs, that are often considered in the literature, see e.g. [20, 22, 75, 25, 26].. The Lagrangian of this theory is
@@ 2.2 String ansatz and boundary conditions
The model described above admits classical solutions that are static and uniform in one spatial direction [79]. They are characterised by a pair of field configurations, $Z^{\,\rm str}_{\mu}(x_{1},x_{2})$ and $\Phi^{\rm str}(x_{1},x_{2})$, that have no time ($t$) and $z$ ($x_{3}$) dependence and correspond to a string sourced by the $Z$ and Higgs fields along the $x_{3}$-direction:
@@
The field configurations of Eqs. (2.10) and (2.11) are known as the electroweak strings, or $Z$-strings, and amount to a non-Abelian embedding of the Abrikosov-Nielsen-Olesen (ANO) classical string of the Abelian Higgs model [1, 62]. The index $n\in\mathbb{Z}$ is the magnetic flux of the $Z$-string, and in what follows, we will restrict to an elementary string with $n=1$. Generally, the field configurations described by the ansatz of Eqs. (2.10) and (2.11) are not guaranteed or expected to be stable throughout the model’s parameter space.
@@ 2.3 Topology and the semi-local limit
Although not protected by topology, Refs. [41, 46] showed that such electroweak-like string configurations can be classically stable local minima [^2] A note on terminology: A field configuration is called classically stable if it is a local minimum of the energy functional $E$. In particular, there should be no negative or zero modes when expanding $E$ around a classically stable field configuration. There can, and in general will exist other local minima with lower values of $E$ separated by finite or infinite energy barriers from the classical solution. When the barrier is finite, the classical string can decay via quantum tunnelling into the true minimum. In this case the classical string is metastable and, in favourable conditions, its quantum decay rate can be computed semiclassically [64] using a version of Coleman’s false vacuum decay approach [27]. of the energy per unit length in a restricted region of parameter space when the model becomes semi-local:
@@
We demonstrate this stability in the semi-local limit, where the $SU(2)$ sector of the gauge group $G$ becomes a global symmetry, while the $U(1)$ factor remains gauged. The resulting semi-local model then possesses a topologically conserved, integer-valued magnetic flux [77, 65] (see also sec. IID.2 in Ref. [64] for a concise summary of the topological argument), which corresponds to the quantum number $n\in\mathbb{Z}$ appearing in the string field configurations Eqs. (2.10) and (2.11). As mentioned, in what follows, we will restrict to the elementary flux sector with $n=1$. Stability is determined by an energy comparison between two configurations: the localised $Z$-string, where energy is confined in a flux tube of tension $\mu_{\rm str}$, and a delocalised texture, where the same flux spreads across a broad scalar field configuration of tension $\mu_{\infty}$:
@@ 3 Tension Parameters and Classical Stability
This section computes the two dimensionless tension coefficients that control the $Z$-string’s stability and decay rate. In Section 3.1 we derive $\alpha_{\infty}$ (delocalised flux, sigma-model limit); in Section 3.2 we compute the string tension $\alpha_{\rm str}(\beta)$ numerically as a function of the mass ratio $\beta$.
@@ 3.1 Computing the $\alpha_{\infty}$ tension coefficient
To compute $\alpha_{\infty}$, we evaluate the energy of static two-dimensional configurations carrying one unit of magnetic flux, uniformly delocalised over a disc of radius $R$ and taking the $R\to\infty$ limit. Following Ref. [64] (sec. VIIIA), we can make two simplifying assumptions. For these infinitely-spread finite energy configurations, the Higgs field takes values in the vacuum manifold everywhere,
@@
where $g(\Phi)_{\alpha\beta}$ is the metric tensor of the $CP^{1}$ theory in affine coordinates, given by the Fubini-Study metric, see $e.g.$ [31, 42],
@@
The real-valued fields $\varphi^{a}$ defined in Eq. (3.6) satisfy the relation $(\varphi^{1})^{2}+(\varphi^{2})^{2}+(\varphi^{3})^{2}=\eta^{2}/2$ and conveniently parameterise the vacuum manifold of our original Abelian Higgs model,
@@
Here $E$ is the transverse energy, $i.e.$ the 3+13{+}1D string tension (energy per unit length). The model of Eq. (3.9) has topological solitons [63, 70] with topological charges $n\in\pi_{2}(S^{2})\sim\mathbb{Z}$. The single soliton solution is given by,
@@
More detail and a derivation of the soliton solution above can also be found in Sec. 7 of the Ref. [70]. In the sigma model of Eq. (3.9), the soliton size parameter $r_{0}$ is a flat direction. But the sigma model soliton provides an accurate approximation to the large-size solution of our $SU(2)\times U(1)$ semi-local model only for large scale sizes. Hence, we can identify $r_{0}$ with $R$ when we take the latter to infinity. By equating the two-dimensional soliton mass in Eq. (3.11) with the string tension $\mu_{\infty}$ of the infinite-size magnetic flux configuration of Eq. (2.16), we conclude that
@@ 3.2 Computing the string tension coefficient
The dependence of $\alpha_{\rm str}$ on the mass ratio $\beta$ is displayed in the right panel of Figure 1. As expected from the behaviour of the profile functions, $\alpha_{\rm str}$ increases slowly as $\beta$ increases: a heavier Higgs field localises the scalar core and slightly raises the overall tension. The role of the $SU(2)$ gauge coupling $g$ is implicit in the choice of dimensionless variables and does not appear in the energy functional [79], or alter the shape of the curves shown in the right panel of Figure 1, although it re-enters when converting back to physical units.
Figure 1: Left: Profile functions P(r~)P($\tilde${r}) and V(r~)V($\tilde${r}) for the classical string configuration. The functions are plotted for $\beta=0.02,\,0.2$ and 1. These functions are independent of $g$. Right: The blue line shows the dependence of $\alpha_{\text{str}}$ on $\log\beta$, for the range $2<\log_{10}\beta<0$, where the string solution can be shown to be metastable. Note also that $\alpha_{\text{str}}$ depends on $\beta$ only, and has no further dependence on $g$. The grey dotted line shows the analytical approximation for $\alpha_{\rm str}$ valid in the small-$\beta$ limit and the red line $\alpha_{\rm str}=\pi$.
@@
We observe that for $\beta=1$ the string coupling satisfies $\alpha_{\rm str}=\pi$, coinciding with the value for $\alpha_{\infty}$ and saturating the Bogomolnyi bound [17] for the energy. In addition, in the $\beta\to 0$ limit the ANO string tension has been obtained analytically in Ref. [85],
@@ 4 Bounce Calculation
When we turn on $g\neq 0$ the $SU(2)$ group is gauged and the string magnetic flux is no longer protected by topology as it can ‘unwind’ by exciting non-Abelian gauge field degrees of freedom. The string is no longer absolutely stable but there is a region in the parameter space where it is still a local minimum now separated by a finite energy barrier from minima with lower energy. This classical stability region of the string was established in Ref. [46], and is shown in Figure 3 as the area to the right of the black contour on the $(\beta,\sin^{2}\theta_{w})$ plane. The classical string decays by quantum tunnelling and in this section, we will derive the bounce action which determines the semiclassical exponent of the string decay rate following the thin-defect aka thin-wall approach of [64].
@@
can terminate, while the magnetic flux of the $B_{\mu}$ component is converted to the magnetic flux of the massless photon field $A_{\mu}$. The end points of the $Z$-string are electromagnetic monopoles of Nambu [61, 49], [^3] These configurations exist only as string endpoints, i.e. they are always confined by the string and do not appear as free classical solutions. In practice they are ’t Hooft–Polyakov-type monopole configurations composed of the $W_{\mu}^{a}$ gauge field and the adjoint Higgs field $n^{a}=\Phi^{\dagger}\tau^{a}\Phi/(\Phi^{\dagger}\Phi)$. This composite scalar field configuration is not invertible for $\Phi(x)=0$, which is precisely the point where the $Z$-string is attached. they absorb the string’s $Z$-flux and continuously convert it into $A$-flux,
@@
and considering the hierarchy of scales shown in the left panel of Figure 2. Summing these three contributions gives the total energy contained within the monopole core [64]:
@@
where again we expanded in $g^{2}/g^{\prime 2}\ll 1$ [^4] We note that in [64], there is an error in the calculation of $R_{\rm core}^{2}$ by a factor of 2.. Eq. (4.7) is an analytic estimate rather than the result of a full numerical solution of the coupled non-linear field equations for the endpoint configuration in the $SU(2)\times U(1)$ model. In the semilocal limit this expression for the monopole mass is exact in the sense that it provides the leading singularity in the $g\to 0$ limit. More generally, there will be additional contributions on the right hand side of Eq. (4.7) suppressed by extra powers of $g^{2}$. A more precise determination of such corrections is an interesting direction for future work. For now we take the monopole mass to be given by Eq. (4.7).
To determine the bounce action, we follow the Euclidean picture shown in the right panel of Figure 2, reproducing the monopole–antimonopole nucleation process on aă string as described Ref. [64]. A string sweeps out a two‑dimensional worldsheet with one timelike and one spacelike coordinate. Performing a Wick rotation converts the timelike direction into another spacelike one, converting the Lorentzian worldheet into a Euclidean surface. Quantum fluctuations may then punch a circular hole of radius $R$ in the worldsheet and the rim of this hole is the closed worldline of the nucleated monopole–antimonopole pair. In the thin‑wall limit, where the string thickness and the monopole core size are negligible compared with the radius of the critical bubble $R$, the Euclidean action simplifies to
@@
The result is valid as long as all scales remain much smaller than $R$. The monopole mass in this expression corresponds to the energy contained within the monopole core, obtained in Eq. (4.7), and combining these two equations we find, reproducing the result found in Ref. [64],